# Functions of bounded Musielak-Orlicz-type deformation and anisotropic Total Generalized Variation for image-denoising problems

### v3

@@ 4.1 Basic properties of $TGV^{\varphi,2}_{\alpha}(u)$
(i). Thanks to the definition of the functional $TGV^{\varphi,2}_{\alpha}$ given in (1.1), it is obvious to see that the two conditions (N2) and (N3) in Definition 2.1 hold, implying that $TGV^{\varphi,2}_{\alpha}$ is a seminorm. Note that $BGV^{\varphi,2}_{\alpha}(\Omega)$ is a normed space because $\varphi\in\Phi_{w}(\Omega)$ and satisfies (A0), so $\|\cdot\|_{\varphi}$ is a norm, see [29, Theorem 3.7.13].
@@ 4.2 Dual formulation
again for every $w\in{\mathcal{D}}^{\prime}(\Omega;\mathbb{R}^{n})$. Thus, to summarize, by classical properties of $BD$ and $BV$, see [45, Chapter II, Theorem 2.3] and [4], we have that the following characterization holds:
@@
To conclude the proof we observe that if $u\in BV^{\varphi}(\Omega)$, then by [29, Theorem 3.4.6] the above infimum problem can be considered on the subspace of $BD(\Omega)$ where additionally $\|w\|_{\varphi^{**}}<+\infty$, namely $BD^{\varphi}(\Omega)$, see Remark 3.8. ∎
@@ 4.3 Existence and stability of solutions to the minimization problem
Since $\psi_{\varepsilon}$ is a valid test function and $\varphi^{\prime}(\cdot,|g|)<+\infty$ a.e. in $\Omega$, the last equality, together with “Young’s equality”, see [40, Lemma 1.7.3(i)], implies that

### v4

@@ 4.1 Basic properties of $TGV^{\varphi,2}_{\alpha}(u)$
(i). Thanks to the definition of the functional $TGV^{\varphi,2}_{\alpha}$ given in (1.1), it is obvious to see that the two conditions (N2) and (N3) in Definition 2.1 hold, implying that $TGV^{\varphi,2}_{\alpha}$ is a seminorm. Note that $BGV^{\varphi,2}_{\alpha}(\Omega)$ is a normed space because $\varphi\in\Phi_{w}(\Omega)$ and satisfies (A0), so $\|\cdot\|_{\varphi}$ is a norm, see [30, Theorem 3.7.13].
@@ 4.2 Dual formulation
again for every $w\in{\mathcal{D}}^{\prime}(\Omega;\mathbb{R}^{n})$. Thus, to summarize, by classical properties of $BD$ and $BV$, see [46, Chapter II, Theorem 2.3] and [4], we have that the following characterization holds:
@@
To conclude the proof we observe that if $u\in BV^{\varphi}(\Omega)$, then by [30, Theorem 3.4.6] the above infimum problem can be considered on the subspace of $BD(\Omega)$ where additionally $\|w\|_{\varphi^{**}}<+\infty$, namely $BD^{\varphi}(\Omega)$, see Remark 3.8. ∎
@@ 4.3 Existence and stability of solutions to the minimization problem
Since $\psi_{\varepsilon}$ is a valid test function and $\varphi^{\prime}(\cdot,|g|)<+\infty$ a.e. in $\Omega$, the last equality, together with “Young’s equality”, see [41, Lemma 1.7.3(i)], implies that